OpenStudy (anonymous):

True or False The y-value for y = cos(x) when x = -90° is 0.

OpenStudy (johnweldon1993):

Hint* Cos is an odd function...so \(\large \cos(-x) = \cos(x)\) We know \(\large \cos(90) = 0\) so \(\large \cos(-90) = 0\) So true

OpenStudy (anonymous):

thanks